# FERA: Foil Fencing Referee Assistant Using Pose-Based Multi-Label Move Recognition and Rule Reasoning

### v2

@@ Abstract
The sport of fencing, like many other sports, faces challenges in refereeing: subjective calls, human errors, bias, and limited availability in practice environments. We present FERA (FEncing Referee Assistant), a prototype AI referee for foil fencing which integrates pose-based multi-label action recognition and rule-based reasoning. FERA extracts 2D joint positions from video, normalizes them, computes a 101-dimensional kinematic feature set, and applies a Transformer for multi-label move and blade classification. To determine priority and scoring, FERA applies a distilled language model with encoded right-of-way rules, producing both a decision and an explanation for each exchange. With limited hand-labeled data, a 5-fold cross-validation achieves an average macro-F1 score of 0.549, outperforming multiple baselines, including a Temporal Convolutional Network (TCN), BiLSTM, and a vanilla Transformer. While not ready for deployment, these results demonstrate a promising path towards automated referee assistance in foil fencing and new opportunities for AI applications, such as coaching in the field of fencing.
@@ 1 Introduction
Refereeing plays a fundamental role in competitive sports, ensuring fairness and consistency. In fencing, the referee’s job is often more complex than in other sports: they must evaluate blade interactions, interpret right-of-way, and consider who initiated the attack, all at very high speed. Even at official competitions with trained referees, disputes and inconsistent calls still arise due to human error. In club settings, the problem is worse: bouts are often refereed by fellow members or proceed without a referee at all.
The task at hand, however, is more challenging than it first appears. AI refereeing requires robust recognition of temporally dependent actions under occlusion and high speed. Existing fencing action datasets are narrow in scope, covering only a few footwork moves from individual fencers, often under fixed camera angles. As a result, they fail to generalize to real bouts involving two fencers, interactions, and diverse viewing conditions. To overcome this limitation, we created a new dataset and labeling protocol.
Together with this dataset, we introduce FERA (FEncing Referee Assistant), a four-stage pipeline that: (i) 2D pose extraction, (ii) kinematic and temporal feature computation, (iii) a Transformer encoder for multi-label move and blade classification (FERA-MDT), and (iv) a distilled language model that applies encoded right-of-way rules for final decisions (FERA-LM). FERA is trained on a new dataset built to overcome the limitations of prior work. We next situate FERA within the context of related research on fencing action recognition and pose-based tracking, before detailing our dataset and methodology.
@@ 2.1 AI in fencing
AI has been explored in fencing, but existing approaches face notable limitations. FenceNet and related models report over 85% accuracy on the Fencing Footwork Dataset [25, 14], which contains 652 videos from 10 fencers annotated with actions such as step forward, step backward, rapid lunge, incremental-speed lunge, lunge with waiting, and jumping-sliding lunge. While valuable for modeling footwork, the dataset is narrow: it omits many actions, particularly blade-related ones, that are essential for referee decisions in foil. All recordings come from a single environment with fixed, straight-on angles, limiting robustness in real-world conditions. Each video shows only one fencer, without opponent exchanges. As a result, the clips represent best-case scenarios, lacking the reflexive responses and corrections that occur in actual bouts.
Although such conditions simplify training and yield high reported accuracy, models trained on this data cannot generalize to full foil bouts with diverse speeds, movements, and camera perspectives. Our dataset and labeling protocol address these shortcomings, albeit at the cost of increased modeling difficulty. To meet this challenge, FERA builds on pose-based analysis and robust multi-person tracking to extract reliable motion cues in real bouts.
@@ 2.2 Pose Estimation and Tracking
Human pose estimation is central to analyzing athletic performance because it converts raw video into structured joint representations, reducing noise. While direct video pre-processing can yield results, pose estimation has proven more effective for extracting relevant motion information. Early work, such as OpenPose, introduced Part Affinity Fields to detect body parts and associate them with individuals in crowded scenes [4], enabling multi-person keypoints and downstream sport applications. Later top-down approaches improved accuracy by cropping per-person bounding boxes before extraction [22]. More recently, vision transformers (e.g., ViTPose) have advanced accuracy further through large-scale pretraining and attention [23]. In this work, we use MMPose [15], which provides multiple pose-extraction frameworks. Specifically, FERA utilizes RTMPose [9] for pose estimation and RTMDet [13] for bounding box detection. RTMPose is lightweight, real-time, and maintains competitive accuracy.
Accurate per-frame keypoints are only part of the solution. Tracking identities and poses over time, especially under occlusion from referees, spectators, or when fencers cross paths, is challenging. Multi-object trackers such as DeepSORT [20] link detections using motion cues and can re-connect after short occlusions. Newer trackers like ByteTrack [24] retain all detections, including low-confidence ones, to avoid losing objects during fast, blurry frames. FERA instead employs a custom tracker tailored for fencing, combining a distance function with Norfair [19, 18].
A further challenge is reliably detecting the fencers themselves. No fencing-specific detector currently exists to distinguish the two active fencers from referees or spectators. Prior work assumes either single-subject lab recordings or generic multi-person tracking, which cannot guarantee correct selection. Ignoring this issue undermines the pipeline: the chance of selecting the wrong subject among many is high. Even filtering for the most confident poses often selects the referee, who is more static, rather than the fencer performing a fast action.
3 Data
To overcome the limitations of prior single-fencer, single-angle datasets, we constructed a dataset of fencing bout clips designed to capture two-fencer exchanges under varied camera angles. The dataset is derived from professional match videos (720p, 25 fps) from three Grand Prix competitions involving approximately 130 competitors. Each match video is segmented into the final seven seconds preceding a touch. Referee decisions are inferred by analyzing the score: if the score changes before another touch occurs, a point is awarded. By aligning these score changes with detected hit events and applying the rules, we obtain referee decisions associated with the clips. This process yields a large set of segments paired with referee outcomes, which form the basis of the rule-reasoning dataset. Each clip is then manually annotated frame by frame to identify the moves performed, providing the data required for FERA-MDT. Details of hit detection, score extraction, and label assignment are given in Section 3.1.
@@ 3.1 Dataset specifics and labeling
Many fencing videos display static overlays for scoring and hit lights. We manually locate the coordinates where the hit lights appear and the score digits are shown. In these regions, reference images are compared to video frames; if the color similarity exceeds tuned thresholds, a hit is detected. For score extraction, number recognition is applied to the identified score zones.
With clips and touch times determined, actions are annotated frame by frame using a custom graphical interface (Figure 1). Each clip is segmented into moves and labeled with one or more of the following actions: step forward, step backward, half step forward, half step backward, lunge, flèche, wait, parry, beat, counterattack, fake, and hit. Each segment is also assigned a blade position from {4, 6, 7, 8, other}. In cases where moves overlap (e.g., beat + step forward), the interval is marked by the endpoints of the longer action, and both labels are retained.
Note to fencers: the hit label may appear arbitrary, but it is simple: it indicates the fencer extended their arm with the intention to score. It does not guarantee that contact was successful.
Figure 1: Custom graphical labeler used to annotate moves and blade positions frame by frame.
In Figure 1, the fencer poses are already visualized in the video. This is done so that, in case the fencer selector fails and chooses an incorrect person, we omit the entire clip to prevent introducing incorrect data.
@@ 3.2 Data Distribution
In total, the dataset contains 1,734 videos, covering 2,386 move instances, with an average of 1.38 moves per video. The moves are skewed, with the most common actions being step forward (28.8%), hit (19.9%), and step backward (14.6%), while rare moves such as flèche account for less than 1% (Table 1(a)).
@@
Blade positions are highly imbalanced: position 6 accounts for 64.2% of annotations, followed by position 8 (23.6%), while positions 4, 7, and other together represent only 12.2% (Table 1(b)). This skew is difficult to mitigate, since fencers typically favor one or two blade positions that they switch between during a match. As most competitors rely on positions 6 and 8 far more than 4 or 7, instances combining a rare move with a rare blade position are exceedingly uncommon.
Data ethics and usage. All videos used in this study are publicly available on YouTube. The dataset is used strictly for academic research under fair use and is not intended for commercial purposes. All annotations were created manually by fencing experts. Because no direct human subjects were involved, informed consent was not required. All released data are fully anonymized.
@@ 4.1 FERA pipeline overview
The FERA pipeline converts raw video into structured referee decisions through four stages: (i) pose estimation, (ii) feature engineering, (iii) Transformer-based move and blade detection (FERA-MDT), and (iv) language-model reasoning (FERA-LM). This overview follows that flow and introduces the dynamic windowing mechanism used to determine which frames are passed to FERA-MDT.
FERA begins by extracting 2D skeleton poses from video, which are then normalized and augmented with the kinematic features described in Section 4.3. Frame sequences are next fed into FERA-MDT and evaluated for detection confidence. If confidence is insufficient, the window expands incrementally by one frame, up to a maximum of forty. If the limit is reached without a confident detection, the starting frame shifts forward, and the process repeats. FERA continues until a confident detection is made. For half-steps, the system checks for a subsequent full step before finalizing the label. This dynamic window accommodates a wide variation in move length; static windows tend to overlap actions and create boundary errors. Once all moves are identified, the predicted actions and blade positions for both fencers are combined with encoded priority rules and passed to FERA-LM, which produces a concise decision and explanation. The following sections describe each stage in detail, beginning with fencer selection and tracking.
Figure 2: Pipeline visualization.
@@ 4.2 Pose Estimation
The first stage of FERA requires accurately identifying the two active fencers and tracking their skeletons across frames, even under occlusion. This process consists of three components: fencer filtering, tracking, and joint extraction.
@@
At the start of each clip, a grace period is used to search for candidate fencers. Candidates must occupy at least 3% of the frame area, exceed a confidence threshold (>0.20>0.20), and not touch the bottom edge (referees are often visible only from the torso upward). Each candidate is then scored with a criterion tuned to select fencers, prioritizing size and penalizing positions that are too high or low in frame.
@@
Tracking is implemented using Norfair and a custom distance function that combines intersection-over-union (IoU) and centroid distance [19, 18]. If an ID is lost, the system attempts to reassign it by matching new detections to the last known bounding box via IoU overlap. Pose similarity scoring further prevents switches to background individuals. A light exponential moving average (EMA) smooths the joint trajectories, retaining valid positions when detections are missing and reducing jitter.
@@
Bounding boxes and skeletons are obtained using MMDetection and MMPose [5, 15]. Specifically, we employ rtmpose-m_8xb256-420e_coco-256x192 for pose extraction and rtmdet_tiny_8xb32-300e_coco for detection, selected as the fastest available models within RTMPose [9] and RTMDet [13] while maintaining competitive accuracy.
@@ 4.3 Feature Engineering
The second stage of FERA involves feature engineering. While raw joints are informative, engineered features capture temporal and geometric dynamics that joints alone cannot convey.
After stage one, 17 joints are available. Because head joints are unreliable due to masks, we discard five of them, leaving 12 body joints. To remove background dependence, we compute a scale- and translation-invariant representation by centering each frame at the pelvis and normalizing by average torso length:
@@
From these normalized joints, we derive additional features:
Pairwise distances:
$d_{ab,t}=\|\tilde{\mathbf{p}}_{a,t}-\tilde{\mathbf{p}}_{b,t}\|,$
for key pairs such as hand-to-hand, foot-to-foot, and shoulder/hip widths.
Joint angles: at elbows and knees. For three joints aabbcc,
@@
Torso orientation: represented as $\sin$ and $\cos$ of the angle between shoulder and hip centers.
Arm extension: magnitude and normalized direction of shoulder–wrist vectors.
Center of mass (CoM): average of all valid joints.
To capture dynamics, temporal derivatives are also included:
@@
Altogether, each frame is represented by a 101-dimensional vector: 24 joint coordinates, 2 CoM features, 11 distances, 4 angles, 2 torso orientation features, 6 arm-extension features, 24 velocities, 24 accelerations, and 4 CoM velocity/acceleration features. These descriptors serve as the input tokens to FERA-MDT.

### v3

# FERA: Bridging the Semantic Gap in Foil Fencing via Kinematic Pose Recognition and Explainable Rule Reasoning

@@ Abstract
Foil fencing presents a unique multimedia challenge: actions are extremely fast, interactions are subtle, and the final right-of-way decision relies on complex linguistic rules applied to visual data. We present FERA (FEncing Referee Assistant), a pose-based framework that bridges the gap between pixel-level perception and semantic rule reasoning. From monocular video, FERA extracts 2D poses, converts them into a compact 101-dimensional kinematic representation, and applies an encoder-only Transformer (FERA-MDT) to predict multi-label footwork and blade actions. Crucially, these predictions serve as semantic tokens for a downstream language model (FERA-LM) to generate explainable verdicts. Training treats left and right fencers symmetrically by creating two single-fencer sequences per clip. At inference, FERA-MDT uses dynamic temporal windowing to handle variable-length clips, and paired predictions are passed to FERA-LM, which applies encoded right-of-way rules to generate prototype decisions and short explanations. On 1,734 clips (2,386 move instances) from professional bouts, FERA-MDT reaches a macro-F1 of 0.549 under 5-fold cross-validation, outperforming a BiLSTM, TCN, and baseline Transformer. Furthermore, we demonstrate that our structured outputs enable a language model to perform logical rule reasoning, effectively decoupling visual perception from rule application. FERA provides the first dataset and benchmark for this cross-modal action understanding task.
@@ 1 Introduction
In fencing, referees must evaluate blade interactions, interpret right-of-way, and determine who initiated or interrupted the attack at high speed. Even at official competitions, disputes arise; in clubs, bouts are often refereed by fellow members or proceed without a referee at all.
Automated assistance in fencing is not merely a computer vision problem; it is a multimodal translation challenge. The system must bridge the semantic gap: translating raw high-speed video pixels into geometric pose data, then into discrete semantic actions (e.g., Parry, Lunge), and finally into a linguistic explanation of the rule outcome. While electronic scoring systems indicate when a touch occurs, they cannot explain why a point was awarded.
Problem definition.
Given a short monocular video clip (approximately 1–3 seconds) centered on a foil exchange, the goal is to:
  1. 1.
    track both fencers,
  2. 2.
    for each fencer, predict a set of footwork and blade actions (multi-label) and a single blade-line position, and
  3. 3.
    consume these per-fencer predictions in a reasoning module that outputs a right-of-way decision and short explanation.
This separates perception (pose-based multi-label action and blade recognition) from reasoning (mapping paired move sequences to referee decisions), allowing each to be evaluated and improved independently.
In our experiments we adopt a trimmed-clip setting for supervised training and cross-validation, where each example corresponds to a manually labeled single-fencer exchange. At inference on longer or untrimmed bout streams, we instead run FERA-MDT with a dynamic temporal windowing scheme over per-fencer pose tracks, so the model does not assume that exchange boundaries or move start/end frames are known in advance.
To support this task, we created a new dataset and labeling protocol. Each clip is drawn from a professional bout and assigned a unique identifier. For recognition, we adopt a canonical single-fencer view: the model always sees a fencer on the left side of the piste. Practically, we annotate the left fencer in the original video, then create a horizontally flipped version of the clip and annotate the original right fencer, who now appears on the left. This yields up to two single-fencer sequences per exchange (one per side), linked by the same clip ID, and lets FERA-MDT train on one skeleton at a time while learning from both fencers.
FERA framework.
FERA (FEncing Referee Assistant) is a four-stage pipeline:
  1. 1.
    extract 2D poses for both fencers from monocular video,
  2. 2.
    compute kinematic and temporal features from each fencer’s pose sequence,
  3. 3.
    apply an encoder-only Transformer for per-fencer multi-label move and blade classification (FERA-MDT), and
  4. 4.
    consume paired left/right predictions in a distilled language model that applies encoded right-of-way rules (FERA-LM).
FERA is trained and evaluated on our dataset, the FERA-Dataset. In this paper we focus on the motion recognition and calibration aspects of FERA-MDT and provide a qualitative demonstration of language-based reasoning. The dynamic temporal windowing stage allows the same model trained on trimmed single-fencer segments to be deployed on arbitrary-length per-fencer pose tracks.
Contributions.
This paper:
  • introduces the FERA-Dataset, a foil fencing benchmark built from two-fencer exchanges, with frame-level annotations of footwork, blade actions, and blade-line positions represented as single-fencer sequences in a canonical left-facing view,
  • proposes FERA-MDT, an encoder-only Transformer for pose-based multi-label move and blade recognition, with analysis of calibration and error patterns against strong baselines (BiLSTM, TCN, and a vanilla Transformer),
  • demonstrates a decoupled vision-to-language reasoning pipeline (FERA-LM), showing that structured action recognition outputs can serve as semantic tokens for rule-based explanation generation, and
  • releases annotations, derived pose features, and baseline code to support research on pose-based motion understanding and AI-assisted sports officiating.
@@ 2.1 AI in fencing
FenceNet and related models report over 85% accuracy on the Fencing Footwork Dataset [28, 17], which contains 652 videos from 10 fencers annotated with footwork actions such as step forward, step backward, and several lunge variants. While useful for footwork modeling, the dataset omits blade actions that are essential for foil refereeing and uses a single recording environment with fixed, straight-on angles. Each video shows only one fencer, without opponent exchanges or reflexive responses.
These conditions simplify training and boost reported accuracy but do not reflect full bouts with two interacting fencers, diverse speeds, and varied camera perspectives. Our dataset and labeling protocol address these limitations by (i) accounting for two interacting fencers, (ii) including both footwork and blade actions, and (iii) using real competition footage with varied viewpoints, at the cost of increased modeling difficulty.
@@ 2.2 Pose estimation and tracking
Human pose estimation converts raw video into structured joint representations for downstream analysis. OpenPose introduced Part Affinity Fields to detect body parts and group them into individuals [4]. Later top-down methods improved accuracy by cropping per-person bounding boxes before extraction [25], and vision transformers such as ViTPose further improved accuracy through large-scale pretraining [26]. We use MMPose [18] with RTMPose [12] for pose estimation and RTMDet [16] for bounding box detection; RTMPose is lightweight and real-time while maintaining competitive accuracy.
Tracking identities and poses over time, especially under occlusions or when fencers cross paths, is challenging. Multi-object trackers such as DeepSORT [23] and ByteTrack [27] link detections over time, with ByteTrack retaining low-confidence detections to avoid losing objects during fast, blurry frames. FERA instead uses a custom tracker tailored for fencing, combining a distance function with Norfair [22, 21] to maintain two consistent fencer tracks per clip.
2.3 AI-assisted refereeing in sports
Beyond fencing, AI-assisted refereeing and explainable decision systems have been explored in sports such as soccer. Recent work on Video Assistant Referee Systems (VARS) develops multi-view pipelines for foul detection and offside decisions and uses large language models for human-interpretable explanations [9, 10, 6]. Benchmarks such as SoccerNet explicitly target temporal localization of fouls, multi-modal reasoning, and explanation quality in realistic broadcast footage.
FERA extends this line of work to foil fencing, where exchanges are shorter, motion is faster and more linear, and right-of-way depends on subtle blade timing rather than continuous physical contact. We focus on pose-based single-fencer action recognition and its calibration; language-based reasoning over recognized actions builds directly on recent VARS-style approaches.
3 FERA-Dataset
We constructed the FERA-Dataset to capture realistic two-fencer exchanges under varied camera angles. It is derived from professional match videos (720p, 25 fps) from three Grand Prix competitions involving approximately 130 competitors. Each match video is segmented into the final seven seconds preceding a touch. Referee decisions are inferred by analyzing the score: if the score changes before another touch occurs, a point is awarded. Aligning score changes with detected hit events yields clips paired with referee outcomes for the rule-reasoning subset. Each clip is then manually annotated to identify the moves performed by each fencer.
@@ 3.1 Dataset specifics and labeling
Many fencing videos display static overlays for scoring and hit lights. We manually locate the hit-light and score regions and compare reference images to video frames; if the color similarity exceeds tuned thresholds, a hit is detected, and digit recognition recovers the score. These signals are used only to align clips with referee decisions; all move and blade annotations are manual.
Actions are annotated frame by frame using a custom interface (Figure 1). Each clip has a unique identifier. We adopt a canonical view in which the fencer of interest always appears on the left:
  • In the original clip, annotators label the fencer on the left side of the piste.
  • In a flipped view, the clip is mirrored horizontally and annotators label the original right fencer, now on the left.
In both passes, the labeling scheme is identical. For the visible (left) fencer, each continuous segment is assigned one or more of:
step forward, step backward, half step forward, half step backward, lunge, flèche, wait, parry, beat, counterattack, fake, hit
and one blade position from {4, 6, 7, 8, other}. In cases where moves overlap (e.g., beat + step forward), the interval is marked by the endpoints of the longer action, and both labels are retained.
Note to fencers: the hit label indicates that the fencer extended their arm with the intention to score and does not guarantee that contact was successful.
Figure 1: Custom labeler for per-fencer moves and blade positions. Visualized pose estimates help identify labeling or tracking failures; clips with unreliable fencer selection are discarded.
Each annotated example can be viewed as a single-fencer sequence of time-stamped records:
$\big\{(t_{k},\mathcal{Y}_{k},b_{k})\big\}_{k=1}^{K},$
where $\mathcal{Y}_{k}\subseteq\{1,\ldots,12\}$ is the set of active move labels for the (left-view) fencer at interval $k$, and $b_{k}\in\{4,6,7,8,\text{other}\}$ is the corresponding blade position. When both sides are annotated, a clip with identifier $c$ yields two sequences, $(c,\text{left})$ and $(c,\text{right})$, which share the same time axis but are stored and trained as separate single-fencer examples.
FERA-MDT is trained purely on these single-fencer sequences in the canonical left view. At inference, we extract pose tracks for each fencer, mirror the right-side track if needed, apply FERA-MDT independently using dynamic temporal windowing (Section 4), and then realign the resulting move and blade sequences on the shared time axis before passing them to FERA-LM.
@@ 3.2 Data distribution
We use 1,734 clips and 2,386 annotated move instances across all single-fencer sequences. On average, each clip contains 1.38 moves. The moves are skewed, with step forward (28.8%), hit (19.9%), and step backward (14.6%) most common, while rare moves such as flèche account for less than 1% (Table 1(a)).
@@
Blade positions are highly imbalanced: position 6 accounts for 64.2% of annotations, followed by 8 (23.6%), while 4, 7, and other together represent 12.2%. This reflects typical foil tactics, where most fencers rely on one or two blade lines, making rare move/blade combinations extremely uncommon.
Data ethics and availability. All source videos are publicly available competition recordings. We work only with broadcast footage and do not collect additional identifying information. All annotations were created manually by fencing experts. Released data are anonymized and consist of derived pose sequences and labels. We provide a research-access package containing pose sequences, kinematic features, annotations, and training/evaluation code for all baselines; the project page URL is in the supplementary material.
Task formulation. We adopt a trimmed-clip recognition setup for training and evaluation: each example is a short clip centered on an exchange, and the goal is to predict the set of moves and blade position active in that interval for a single fencer. In the deployed pipeline, however, we run FERA-MDT with the dynamic temporal windowing scheme described in Section 4 on arbitrary-length per-fencer pose sequences extracted from bout footage, so clip-level exchange boundaries are not required at test time. Extending FERA to fully annotated action spotting on untrimmed broadcasts—localizing and classifying all actions over entire bouts and evaluating them directly—remains important future work.
@@ 4.1 FERA pipeline overview
The FERA pipeline converts raw video into structured referee inputs and, optionally, decisions through four stages: (i) pose estimation, (ii) feature engineering, (iii) Transformer-based move and blade detection (FERA-MDT), and (iv) language-model reasoning (FERA-LM). FERA operates on one fencer at a time during recognition. For each clip or longer bout segment, pose and feature sequences for the left and right fencers are extracted separately, processed by a shared FERA-MDT model with dynamic temporal windowing, and decoded into per-fencer move sequences and blade positions. Predictions are then aligned on the shared time axis, combined with encoded priority rules, and optionally passed to FERA-LM. Dynamic temporal windowing bridges the gap between training on trimmed single-fencer examples and inference on continuous per-fencer pose streams.
Figure 2: FERA pipeline: pose estimation, feature engineering, move/blade detection (FERA-MDT), and rule reasoning (FERA-LM).
@@ 4.2 Pose estimation
The first stage of FERA identifies the two active fencers and tracks their skeletons across frames, under occlusion and motion. This stage includes fencer filtering, tracking, and joint extraction.
@@
At the start of each clip, we search for candidate fencers during a short grace period. Candidates must occupy at least 3% of the frame area, exceed a confidence threshold (>0.20>0.20), and not touch the bottom edge (referees are often visible only from the torso upward). Each candidate is scored by size and vertical placement to prioritize likely fencers.
@@
Tracking uses Norfair with a custom distance function combining intersection-over-union (IoU) and centroid distance [22, 21]. If an ID is lost, the system reassigns it by matching new detections to the last known bounding box via IoU, with pose similarity scoring to prevent switches to background individuals. A light exponential moving average (EMA) smooths joint trajectories, retaining valid positions when detections are missing and reducing jitter.
@@
Bounding boxes and skeletons are obtained using MMDetection and MMPose [5, 18]. We employ rtmpose-m_8xb256-420e_coco-256x192 for pose extraction and rtmdet_tiny_8xb32-300e_coco for detection, chosen as fast RTMPose [12] and RTMDet [16] models with competitive accuracy.
@@ 4.3 Feature engineering
The second stage derives features from pose sequences. Raw joints are informative, but engineered features capture temporal and geometric structure that joints alone do not. We express all skeletons in a canonical left-facing coordinate system: if the active fencer is originally on the right, their 2D joint coordinates are mirrored horizontally about the frame’s vertical midline. This matches the labeling convention in Section 3.1.
From the 17 available joints we discard five head joints, leaving 12 body joints. To remove background dependence, we compute a scale- and translation-invariant representation by centering each frame at the pelvis and normalizing by average torso length:
@@
From these normalized joints, we derive:
  • Pairwise distances $d_{ab,t}=\|\tilde{\mathbf{p}}_{a,t}-\tilde{\mathbf{p}}_{b,t}\|$ for hand-to-hand, foot-to-foot, and shoulder/hip widths.
  • Joint angles at elbows and knees:
@@
  • Torso orientation as $\sin$ and $\cos$ of the angle between shoulder and hip centers.
  • Arm extension as magnitude and normalized direction of shoulder–wrist vectors.
  • Center of mass (CoM) as the average of all valid joints.
To capture dynamics, we include temporal derivatives:
@@
Each frame is represented by a 101-dimensional vector: 24 joint coordinates, 2 CoM features, 11 distances, 4 angles, 2 torso orientation features, 6 arm-extension features, 24 velocities, 24 accelerations, and 4 CoM velocity/acceleration features. These descriptors serve as input tokens to FERA-MDT.